OpenStudy (anonymous):

Help I will rate and fave your comments!!! 1. 2(x - 3) = 2x a. one solutions b. no solutions c. infinitely many solutions 2.3(y - 3) = 2y - 9 + y a. one solutions b. no solutions c. infinitely many solutions 3. 10x - 2 - 6x = 3x - 2 +x a. one solutions b. no solutions c. infinitely many solutions 4. 4(x +3) +2x = x - 8 a. one solutions b. no solutions c. infinitely many solutions

OpenStudy (anonymous):

Please Help!!

OpenStudy (anonymous):

:3

OpenStudy (arthur_ser):

1. 2(x - 3) = 2x 2x-6=2x -6=0 >no solutions 2. 3(y - 3) = 2y - 9 + y 3y-9=3y-9 0y=0 >infinitely many solutions 3. 10x - 2 - 6x = 3x - 2 +x 4x=4x 0x=0 >infinitely many solutions 4. 4(x +3) +2x = x - 8 4x+12 +2x=x-8 5x=-20 x=-4 >one solutions

OpenStudy (anonymous):

thanks so much!